An airplane travels 280m down the runway before taking off. If it starts from rest, moves with constant acceleration, and becomes airborne in 8.0s, what is it’s speed in m/s when it takes off?



1
Expert's answer
2021-03-25T19:28:57-0400
"d=\dfrac{1}{2}(v_0+v_f)t,""v_f=\dfrac{2d}{t}=\dfrac{2\cdot280\ m}{8.0\ s}=70\ \dfrac{m}{s}."